LET 0>A>45 if 420(tanA+cotA)=84 Then find the value of(116cosA-58sinA)

Grade:10

1 Answers

Rinkoo Gupta
10 years ago
if your question is 420(tanA+cotA)=841
then
420/sinAcosA=841
840/sin2A=841
sin2A=840/841
cos2A=sqrt of (1-sin^22A)=41//841
cosA=21/29
sinA=20/29

hence 116cosA-58sinA=116(21/29)-58(20/29)=44
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty